OpenStudy (khally92):

A large cylindrical soup has a volume of 3456pie cm^2. Find the radius and the height of the can that requires a minimum amount of material to construct it V=pie r^2 h and S= 2 pie r^2 + 2pie r h. OPTIMIZATION PROBLEM

OpenStudy (khally92):

pleaseeeeee.........

OpenStudy (unklerhaukus):

so you want the smallest surface area

OpenStudy (khally92):

yeah

OpenStudy (khally92):

how do i go about it

OpenStudy (unklerhaukus):

\[ S= 2 \pi r^2 + 2\pi r h=2\pi r(r+h)\]

OpenStudy (khally92):

yeah

OpenStudy (unklerhaukus):

\[V=\pi r^2\]

OpenStudy (khally92):

V=pie r^2h

OpenStudy (unklerhaukus):

whops \[V=\pi r^2h\]

OpenStudy (khally92):

yeah

OpenStudy (unklerhaukus):

\[V=3456\pi [\text{cm}]^2\] so can you find \[r^2h=\]

OpenStudy (unklerhaukus):

then find \(h(r)=\)

OpenStudy (khally92):

How?

OpenStudy (unklerhaukus):

did you find r^2h?

OpenStudy (khally92):

r^2h=3456

OpenStudy (unklerhaukus):

now divide by r^2

OpenStudy (khally92):

what divide by r^2

OpenStudy (unklerhaukus):

\[r^2h=3456\] \[h=\frac{3456}{r^2}\]\[h(r)=\frac{3456}{r^2}\]

OpenStudy (khally92):

yeah..

OpenStudy (unklerhaukus):

\[S=2\pi r\left(r+\frac{3456}{r^2}\right)\]

OpenStudy (khally92):

sure am gettin it

OpenStudy (unklerhaukus):

now differentiate \(S\) with respect to \(r\)

OpenStudy (unklerhaukus):

if might be easier to differentiate if you expand the brakets first

OpenStudy (khally92):

i.e 2pie r^2 +6912pie r/r

OpenStudy (unklerhaukus):

\[S=2\pi \left(r^2+\frac{3456}{r}\right)\] \[\frac{\text dS}{\text dr}=\]

OpenStudy (khally92):

how should i find the derivate

OpenStudy (unklerhaukus):

\[y=x^n\]\[\frac{\text dy}{\text dx}=nx^{n-1}\]

OpenStudy (khally92):

wats the derivative of 2pie?

OpenStudy (unklerhaukus):

\[2\pi=2\pi\times1 =2\pi x^0\]

OpenStudy (khally92):

are you saying is constant i tot is -0

OpenStudy (unklerhaukus):

the derivative of a constant is zero yes

OpenStudy (khally92):

so the derivative is 2r+3456r^-2

OpenStudy (unklerhaukus):

the derivative is \[\frac{\text dS}{\text dx}=2\pi\left(2r+\frac{3456}{r^2}\right)\]

OpenStudy (khally92):

you said the derivative to 2pir is =0

OpenStudy (unklerhaukus):

dont forget the chain rule \[(fg)'=f'g+gf'\] when the derivative is zero there is a critical point in the solution curve

OpenStudy (unklerhaukus):

(i mean the product rule not the chain rule ) now. when is \[\left(2r+\frac{3456}{r^2}\right)=0\] solve for r

OpenStudy (khally92):

r=-12

OpenStudy (unklerhaukus):

yeah, how'd you get that ? your right by the way)

OpenStudy (unklerhaukus):

except the radius cant really be negative

OpenStudy (khally92):

i solved for r as you said. yeah plus or negative r. there fore r=12

OpenStudy (unklerhaukus):

ops i made a mistake on the derivative it should be \[\frac{\text dS}{\text dx}=2\pi\left(2r-\frac{3456}{r^2}\right)\]so the critical point is when\[2r-\frac{3456}{r^2}=0\]\[2r=\frac{3456}{r^2}\]\[r^3=\frac{3456}{2}\]\[r=\sqrt[3]{\frac{3456}{2}}=\]

OpenStudy (khally92):

ok r=12 how we solve fot h now.

OpenStudy (unklerhaukus):

yes!

OpenStudy (khally92):

how should i solve for h

OpenStudy (anonymous):

A Mathematica solution is attached.

OpenStudy (unklerhaukus):

\[r=12\] \[h(r)=\frac{3456}{r^2}\] \[h(12)=\frac{3456}{12^2}\] \[h=\frac{3456}{144}=\dots\]

OpenStudy (khally92):

h=24. thanks alot and to robotobey thanks alot too.

OpenStudy (unklerhaukus):

the general case of optimisation of a cylinder is h=2r